# Entanglement dynamics in double Jaynes-Cummings model and intensity-dependent double Jaynes-Cummings model for squeezed coherent thermal states

### v3

@@ 3.1 Wigner function distributions for SCTS
Before, to proceed further in this paper to investigate the effects of different interactions, we discuss the Wigner function $W(\alpha)$ for SCTS in IDDJCM and DJCM. Wigner function distributions are of major importance in quantum optics. $W(\alpha)$ is defined as[74, 75, 76]
@@
Wigner functions $W(\alpha)$ for SCTS in IDDJCM are depicted in Fig. 4. In these plots $\bar{n}_{c}=2.0,\bar{n}_{s}=0.1$ and $\bar{n}_{th}=0.1$ are chosen. It is observed that at $t=0$, $W(\alpha)$ is positive and it a shifted Gaussian; however, at $t=2$ s the peak of $W(\alpha)$ decreases significantly and at some region it becomes negative. The negative peak is almost $30\%$ of the peak at $t=2$ s. After that, the negativity of $W(\alpha)$ decreases and at $t=6$ s, it almost resembles the initial peak at $t=0$ s. As time increases further, it is noticed that negativity of Wigner function increases again at $t=8$ s and again it decreases. So, it is observed that $W(\alpha)$ changing value from positive to negative and vice versa as time progresses. It indicates that the state of radiation oscillating between a classical state and non-classical state as an effect of atom-field interaction in IDDJCM.
@@
Figure 5 represents the Wigner function distributions for SCTS in DJCM. In this case also, $W(\alpha)$ starts from a shifted Gaussian shape at $t=0$s. At $t=2$s it’s positive peak comes down very drastically and in some region it becomes negative (almost $50\%$) of the highest of value. As time increases, the shape of $W(\alpha)$ changes and negative values decreases. However, it is contrasting compared to the case with IDDJCM that $W(\alpha)$ does not become positive again and mimic the initial state. The Wigner function for SCTS shows negative values for a longer period of time in DJCM. It shows that SCTS become more non-classical as a result of atom-field interaction in DJCM as compared to the case in IDDJCM.
@@ 4 Effects of single photon exchange interaction on the entanglement dynamics of the atom-field system
Now we study the effects of single photon exchange between the two cavities on entanglement dynamics. Recently, people have been studying systems in which photons can hop from one atom-cavity system to another. Pandit et. al, [28, 31] have studied the effects of cavity-cavity interaction on the entanglement dynamics of a generalized double Jaynes-Cummings model. The authors have shown that for larger value of $\kappa$, entanglement for the atom-atom subsystem can be protected. They have also investigated the role of $\kappa$ in entanglement transfer between atom-atom and field-field subsystems. In another work, Laha [33] has investigated the effects of beam-splitter which is equivalent to photon exchange interaction on entanglement dynamics in DJCM for radiation states in vacuum, coherent and thermal states. The effects of photon exchange interaction is modeled by the following Hamiltonian:
$\hat{H}_{\text{PE}}=\hat{H}+\kappa\hat{a}^{\dagger}{\hat{a}}+\kappa\hat{b}^{\dagger}\hat{b}$ (42)
@@
$\hat{H^{\prime}}_{\text{PE}}=\hat{H^{\prime}}+\kappa\hat{a}^{\dagger}{\hat{a}}+\kappa\hat{b}^{\dagger}\hat{b}$ (43)
@@
Figure 6 shows the effects of cavity-cavity single photon exchange interaction on the entanglement dynamics for SCTS in IDDJCM. The strength of this cavity-cavity interaction via single photon exchange is characterized by the parameter $\kappa$. To study the effects of $\kappa$, $\bar{n}_{c}=2,\bar{n}_{s}=1$ and $\bar{n}_{th}=0.1$ are taken. From Fig. 6(a) it is observed that for $\kappa=0.1$ i.e., $\kappa<\lambda$, the peak of atom A-atom B entanglement pulses decreases with time (described by blue curve), however, the lengths of ESDs do not change considerably. The field a-field b entanglement which is represented by the black curve shows different behaviour. If we compare Figs. 6(d) and 1(d), it can be seen that ESDs are removed from the dynamics and the amplitude of $N(t)$ increases with time. For both atom A-field a (green curve) and atom A-field b (red curve), the periodic nature of $N(t)$ gets destroyed however the amplitude of $N(t)$ does not change noticeably. For $\kappa=\lambda$ ($\kappa=1$), peaks of $C(t)$ decreases rapidly with time, it is also observed that the length of the first ESD increases however the length of other ESDs decreases at some places. For field-field entanglement, the amplitude of $N(t)$ rises up sharply and then it becomes oscillatory. For the atom A-field a and atom A-field b, all the ESDs are removed from the dynamics though the amplitude of $N(t)$ in these subsystems are small. As we increase $\kappa$ further i.e, $\kappa>\lambda$ ($\kappa=5$), it is observed that ESD in the dynamics of $C(t)$ are removed initially and many smaller peaks appear in the dynamics. From the plot, it is evident that this photon exchange interaction is creating atom-atom entanglement in the system. In the field-field entanglement, $N(t)$ shows reduction in it’s amplitude which can be explained as the increase on number of smaller peaks and disappearance of ESDs in $C(t)$; entanglement is transferred from $N(t)$ to $C(t)$. Also, the frequency of field-field entanglement increases significantly. The atom A-field a entanglement is decreased and it almost becomes zero at some places. $N(t)$ for atom A-field b does not change considerably, however it becomes zero when atom-atom and field-field entanglement are maximum. As we increase the value of $\kappa$ to $10$, we have a situation where $\kappa>>\lambda$ i.e, photon exchange coupling between the cavities is more powerful than coupling between the atoms and radiation field. All the ESDs from the dynamics of $C(t)$ is removed and the value of $C(t)$ increases by a large amount. This increase in atom-atom entanglement makes the field-field entanglement to decrease considerably; entanglement is transferred from field-field subsystem to atom-atom subsystem. $N(t)$ for atom A-field a and atom A-field b are almost zero everywhere except some very little peaks initially. So, we observe that for large photon exchange coupling between the cavities ESDs are removed from atom-atom dynamics and also the amplitude increases significantly. $N(t)$ for field-field entanglement too gets rid of ESDs with the addition of $\kappa$ in the system; however, its amplitude decreases with increasing value of $\kappa$.
@@
Figure 7 shows the effects of photon exchange interaction for SCTS in DJCM. The maximum value of of $N(t)$ for field-field subsystem increases by almost three times as a result of photon exchange interaction. For $\kappa=0.1$, which is a low value compared to the atom-field coupling constant $\lambda$. For this low value of $\kappa$, the lengths of ESD for atom-atom entanglement increase and other smaller peaks disappear. The atom A-field a entanglement comes down significantly and at some places $N(t)$ almost becomes zero. In the case of the atom A-field b entanglement, we observed that the ESDs are removed from the dynamics of $N(t)$. For $\kappa=1.0$, the length of ESD of $C(t)$ becomes larger as small peaks disappear. $N(t)$ for field-field entanglement increases very sharply and then comes down again. It’s maximum value also decreases as compared to the case when $\kappa=0.1$. In atom A-field a subsystem, ESD occurs as a result of increase in $\kappa$. No, considerable change is noticed in atom A-field a entanglement. As $\kappa$ is increased further to $5$, a drastic change is observed in the entanglement dynamics of $C(t)$. We see that all the ESDs are removed from the dynamics and its value also increases significantly at every instant of time. In field-field entanglement, we notice that the maximum value of $N(t)$ increases and the dynamics become very oscillatory. $N(t)$ for atom A-field a and atom A-field b shows initially and after that it stays nonzero with very low amplitude. When $\kappa>>\lambda$ i.e., for $\kappa=10$, $C(t)$ increases more and remains high. We observe interesting pattern in the dynamics of $N(t)$ for field-field entanglement. In this case wave packets of entanglement are formed and amplitude of the wave packets increase significantly. $N(t)$ for atom A-field a and atom A-field b entanglements becomes zero. From this analysis we conclude that photon exchange interaction creates entanglement in the atom-atom and field-field subsystems. However, in the case of IDDJCM, we observe that $C(t)$ increases and $N(t)$ for field-field entanglement decreases; however, in the case of DJCM both $C(t)$ and $N(t)$ for the field-field entanglement increase. In both IDDJCM and DJCM, atom A-field a and atom A-field b entanglements become zero with increasing $\kappa$.
@@ 5 Effects of Kerr-nonlinearity on the entanglement dynamics in IDDJCM
Kerr-nonlinearity affects the dynamics of atom-field interaction and other quantum optical quantities such as the Q-funtion, Wigner function [46]etc. significantly. In ref.[77], the authors have studied the effects of Kerr-nonlinearity and atom-atom coupling on the degree of atom-atom entanglement. In another work, Thabet et. al, [78] have investigated the effects of Kerr-nonlinearity on the mean photon number, Mandel’s $Q$ parameter, entropy squeezing and entanglement dynamics using nonlinear Jaynes-Cummings model. Xi-Wen Hou et. al, [79] have studied the dynamical properties of quantum entanglement in the integrable Jaynes-Cummings model with Kerr-nonlinearity with various Kerr coupling parameters and initial states, where the initial states are prepared by the coherent states placed in the corresponding phase space described in terms of canonical variables.
The entanglement dynamics after adding Kerr-nonlinearity for IDDJCM are presented in Fig. 8 represents The addition of nonlinearity $\chi=0.1$ totally destroys the periodic nature of the entanglement for all the subsystem (blue curve in the plots). ESDs from the dynamics are removed for atom A-field a, atom A-atom b and field a-field b subsystems; however in the case of atom A-atom B, ESDs become longer and peak heights is reduced significantly. As $\chi$ is increased to $0.3$, ESDs get longer in the dynamics of $C(t)$ and also peak which were present now disappear. For other subsystems, amplitude of $N(t)$ for atom A-field a and filed-field entanglement decrease (green curve), however, for atom A-field b subsystem, $N(t)$ increases. For $\chi=0.7$, which is represented by the red curve in the plots, we observe that $C(t)$ starts to revive again making the length of ESDs shorter. In atom A-field a and field a-field b subsystems, amplitude $N(t)$ decreases significantly though there is no ESD present in the system. Like atom-atom entanglement, $N(t)$ increases for atom A-field b subsystem and length of ESDs decreases. Though $N(t)$ decreases at the beginning of the dynamics, it increases with time. So, it can be concluded there is a value $\chi$ for which $C(t)$ becomes minimum and beyond that value it increases again. As we increase the nonlinearity further to $\chi=1.0$, same pattern (black curve) is observed as we see for $\chi=0.7$ i.e., the atom-atom and atom A-field b entanglement increase while field-field and atom A-field a entanglement decreases further.
@@
The effects of $\chi$ on entanglement dynamics of different subsystems for SCTS in DJCM are presented in Fig. 9. For $\chi=0.1$, ESD in $C(t)$ increases considerably since the smaller peaks disappear from the dynamics (Fig. 9(a), blue curve). For the other subsystems dynamics of $N(t)$ does not change considerably for small of $\chi$. As $\chi$ is increased to $0.3$, the amplitudes of $C(t)$ and $N(t)$ increase however, ESDs are not removed from the dynamics. For other entanglements, $N(t)$ for atom A-field a and field a-field b decreases prominently as increasing nonlinearity in the system. However, $N(t)$ for the atom A-field b subsystem increases and ESDs are removed from the dynamics. On further increasing the value of $\chi$ to $0.7$ and $1.0$, $C(t)$ increase very significantly with greater amplitude and ESDs are removed from the dynamics. Similar behaviour is observed for atom A-field b entanglement dynamics. On the other hand, for atom A-field a and field-field entanglement $N(t)$ decreases prominently with these increased values. This shows that addition of $\chi$ in the system, entanglement is transferred in the atom-atom and atom A-field b subsystems from the atom A-field b and field a-field b subsystems.
@@ 6 Effects of spin-spin Ising interaction between the two atoms on entanglement dynamics
where $J_{z}$ is the coupling strength between the two atoms. $J_{z}$ has the unit of energy. Recently people have studied the Ising-type interaction for different configurations. Ghoshal et. al, have investigated the entanglement dynamics of the quenched disordered double Jayens-Cummings model in the presence of Ising-type interaction in the Hamiltonian[31]. In another work, Pandit et. al, have also studied the effects of Ising-type interaction on the entanglement dynamics in the DJCM [28]. In ref.[33] has analyzed the effects of Ising-type interaction on entanglement in DJCM for vacuum, coherent and thermal states of radiation fields. Sadiek et. al,[80] have investigated the time evolution and asymptotic behaviour of entanglement with considering the dipole-dipole and Ising-type interactions in the Hamiltonian.
The effects of Ising interaction on entanglement for IDDJCM are shown in Fig. 10. The plots show very different effects as is observed in DJCM for SCS and G-L states. From Figs. 10(a) and 10(b), it can be seen that for $J_{z}=0.1$ and $0.3$, we cannot distinguish the plots from the plots with $J_{z}=0$ for $C(t)$ and $N(t)$ (see Figs. 1 and 10). However, for atom A-field b entanglement, the amplitude of $N(t)$ comes down and ESDs are also present in the dynamics. In the case of field a-field b entanglement, for the above mentioned values of $J_{z}$, we observe that ESDs are removed from the dynamics and amplitude increases. For $J_{z}=0.7$, small peaks appear along with the large peaks in the dynamics of $C(t)$ (red curve) and also the amplitude decreases. For atom A-field a subsystem, the amplitude of $N(t)$ comes down a little and ESDs are also removed for larger time. In the case of atom A-field b, we do not see any noticeable change in the dynamics. However, for the field a-field b subsystem, the amplitude increases further and no ESD is observed. For $J_{z}=1.0$, small peaks in the dynamics of $C(t)$ appear earlier in the dynamics and its height increases with time while the initial peak decreases. As a result the duration of the ESDs gets smaller. For other subsystems, $N(t)$ follow the dynamics for $J_{z}=0.7$ with decreased peak height for atom A-field a entanglement while increased peak height for atom A-field b and field a-field b subsystems.
@@
The effects of Ising interaction on entanglements in DCJM are depicted in Fig. 11. In this case also, we observe that for small values of $J_{z}$ (0.1 and 0.3), $C(t)$ and $N(t)$ for all the subsystems do not change considerably. Only field-field entanglement increases slightly for $J_{z}=0.3$. For $J_{z}=0.7$, length of ESD for $C(t)$ increase slightly as the small bump disappears from the dynamics. The amplitudes atom-field entanglement do change noticeably, however the length of ESD decreases for atom A-field b entanglement. For $J_{z}=1.0$, two small peaks appear in the dynamics of $C(t)$ which makes the length of ESDs smaller. All the other entanglement dynamics remain almost unchanged. So, we see that Ising interaction is not much effective to remove sudden death of entanglement for atom-atom subsystem and also from the atom A-field b subsystem.

### v4

@@ 3.1 Wigner function distributions for SCTS
Before proceeding further to investigate the effects of different interactions, we discuss the Wigner function $W(\alpha)$ for SCTS in IDDJCM and DJCM. Wigner function distributions are of major importance in quantum optics. $W(\alpha)$ is defined as[74, 75, 76]
@@
Wigner functions $W(\alpha)$ for SCTS in IDDJCM are depicted in Fig. 4. In these plots $\bar{n}_{c}=2.0,\bar{n}_{s}=0.1$ and $\bar{n}_{th}=0.1$ are chosen. It is observed that at $t=0$, $W(\alpha)$ is positive and it is a shifted Gaussian; however, at $t=2$, the peak of $W(\alpha)$ decreases significantly and at some region it becomes negative. The negative peak is almost $30\%$ of the peak at $t=2$ After that, the negativity of $W(\alpha)$ decreases and at $t=6$ , it almost resembles the initial peak at $t=0$. As time increases further, it is noticed that negativity of Wigner function increases again at $t=8$ and again it decreases. So, it is observed that $W(\alpha)$ changing value from positive to negative and vice versa as time progresses. It indicates that the state of radiation oscillating between a classical state and non-classical state as an effect of atom-field interaction in IDDJCM.
@@
Figure 5 represents the Wigner function distributions for SCTS in DJCM. In this case also, $W(\alpha)$ starts from a shifted Gaussian shape at $t=0$. At $t=2$, it is positive peak comes down very drastically and in some region it becomes negative (almost $50\%$) of the highest of value. As time increases, the shape of $W(\alpha)$ changes and negative values decreases. However, it is contrasting compared to the case with IDDJCM that $W(\alpha)$ does not become positive again and mimic the initial state. The Wigner function for SCTS shows negative values for a longer period of time in DJCM. It shows that SCTS become more non-classical as a result of atom-field interaction in DJCM as compared to the case in IDDJCM.
@@ 4 Effects of single photon exchange interaction on the entanglement dynamics of the atom-field system
Now we study the effects of single photon exchange between the two cavities on entanglement dynamics. Recently, people have been studying systems in which photons can hop from one atom-cavity system to another. Pandit et. al., [28, 31] have studied the effects of cavity-cavity interaction on the entanglement dynamics of a generalized double Jaynes-Cummings model. The authors have shown that for larger value of $\kappa$, entanglement for the atom-atom subsystem can be protected. They have also investigated the role of $\kappa$ in entanglement transfer between atom-atom and field-field subsystems. In another work, Laha [33] has investigated the effects of beam-splitter which is equivalent to photon exchange interaction on entanglement dynamics in DJCM for radiation states in vacuum, coherent and thermal states. The effects of photon exchange interaction is modeled by the following Hamiltonian:
$\hat{H}_{\text{PE}}=\hat{H}+\kappa\hat{a}^{\dagger}{\hat{b}}+\kappa\hat{b}^{\dagger}\hat{a}$ (42)
@@
$\hat{H^{\prime}}_{\text{PE}}=\hat{H^{\prime}}+\kappa\hat{a}^{\dagger}{\hat{b}}+\kappa\hat{b}^{\dagger}\hat{a}$ (43)
@@
Figure 6 shows the effects of cavity-cavity single photon exchange interaction on the entanglement dynamics for SCTS in IDDJCM. The strength of this cavity-cavity interaction via single photon exchange is characterized by the parameter $\kappa$. To study the effects of $\kappa$, $\bar{n}_{c}=2,\bar{n}_{s}=1$ and $\bar{n}_{th}=0.1$ are taken. From Fig. 6(a) it is observed that for $\kappa=0.1$, i.e., $\kappa<\lambda$, the peaks of atom A-atom B entanglement pulses decrease with time (described by blue curve), however, the lengths of ESDs do not change considerably. The field a-field b entanglement which is represented by the black curve shows different behaviour. If we compare Figs. 6(d) and 1(d), it can be seen that ESDs are removed from the dynamics and the amplitude of $N(t)$ increases with time. For both atom A-field a (green curve) and atom A-field b (red curve), the periodic nature of $N(t)$ gets destroyed; however, the amplitude of $N(t)$ does not change noticeably. For $\kappa=\lambda$ ($\kappa=1$), peaks of $C(t)$ decrease rapidly with time and also the length of the first ESD increases; however, the length of other ESDs decreases at some places. For field-field entanglement, the amplitude of $N(t)$ rises up sharply and then it becomes oscillatory. For the atom A-field a and atom A-field b, all the ESDs are removed from the dynamics, though the amplitude of $N(t)$ in these subsystems are small. As we increase $\kappa$ further, i.e, $\kappa>\lambda$ ($\kappa=5$), it is observed that ESD in the dynamics of $C(t)$ are removed initially and many smaller peaks appear in the dynamics. From the plot, it is evident that this photon exchange interaction is creating atom-atom entanglement in the system. In the field-field entanglement, $N(t)$ shows reduction in its amplitude which can be explained as the increase on number of smaller peaks and disappearance of ESDs in $C(t)$; entanglement is transferred from $N(t)$ to $C(t)$. Also, the frequency of field-field entanglement increases significantly. The atom A-field a entanglement is decreased and it almost becomes zero at some places. $N(t)$ for atom A-field b does not change considerably; however, it becomes zero when atom-atom and field-field entanglement are maximum. As we increase the value of $\kappa$ to $10$, we have a situation where $\kappa>>\lambda$, i.e, photon exchange coupling between the cavities is more powerful than coupling between the atoms and radiation field. All the ESDs from the dynamics of $C(t)$ is removed and the value of $C(t)$ increases by a large amount. This increase in atom-atom entanglement makes the field-field entanglement to decrease considerably; the entanglement is transferred from field-field subsystem to atom-atom subsystem. $N(t)$ for atom A-field a and atom A-field b are almost zero everywhere except for some very small initial peaks. So, we observe that for large photon exchange coupling between the cavities ESDs are removed from atom-atom dynamics and also the amplitude increases significantly. $N(t)$ for field-field entanglement too gets rid of ESDs with the addition of $\kappa$ in the system; however, its amplitude decreases with increasing value of $\kappa$.
@@
Figure 7 shows the effects of photon exchange interaction for SCTS in DJCM. The maximum value of of $N(t)$ for field-field subsystem increases by almost three times as a result of photon exchange interaction. For $\kappa=0.1$, which is a low value compared to the atom-field coupling constant $\lambda$. For this low value of $\kappa$, the lengths of ESD for atom-atom entanglement increase and other smaller peaks disappear. The atom A-field a entanglement comes down significantly and at some places, $N(t)$ almost becomes zero. In the case of the atom A-field b entanglement, we observed that the ESDs are removed from the dynamics of $N(t)$. For $\kappa=1.0$, the length of ESD of $C(t)$ becomes larger as small peaks disappear. $N(t)$ for field-field entanglement increases very sharply and then comes down again. Its maximum value also decreases as compared to the case when $\kappa=0.1$. In atom A-field a subsystem, ESD occurs as a result of increase in $\kappa$. No, considerable change is noticed in atom A-field a entanglement. As $\kappa$ is increased further to $5$, a drastic change is observed in the entanglement dynamics of $C(t)$. We see that all the ESDs are removed from the dynamics and its value also increases significantly at every instant of time. In field-field entanglement, we notice that the maximum value of $N(t)$ increases and the dynamics becomes very oscillatory. $N(t)$ for atom A-field a and atom A-field b shows ESD initially and after that it stays nonzero with very low amplitude. When $\kappa>>\lambda$, i.e., for $\kappa=10$, $C(t)$ increases more and remains high. We observe interesting pattern in the dynamics of $N(t)$ for field-field entanglement. In this case wave packets of entanglement are formed and amplitude of the wave packets increase significantly. $N(t)$ for atom A-field a and atom A-field b entanglements becomes zero. From this analysis we conclude that photon exchange interaction creates entanglement in the atom-atom and field-field subsystems. However, in the case of IDDJCM, we observe that $C(t)$ increases and $N(t)$ for field-field entanglement decreases; however, in the case of DJCM both $C(t)$ and $N(t)$ for the field-field entanglement increase. In both IDDJCM and DJCM, atom A-field a and atom A-field b entanglements become zero with increasing $\kappa$.
@@ 5 Effects of Kerr-nonlinearity on the entanglement dynamics in IDDJCM
Kerr-nonlinearity affects the dynamics of atom-field interaction and other quantum optical quantities such as Q-funtion, Wigner function [46]etc. significantly. In Ref.[77], the authors have studied the effects of Kerr-nonlinearity and atom-atom coupling on the degree of atom-atom entanglement. In another work, Thabet et. al., [78] have investigated the effects of Kerr-nonlinearity on the mean photon number, Mandel’s $Q$ parameter, entropy squeezing and entanglement dynamics using nonlinear Jaynes-Cummings model. Xi-Wen Hou et. al., [79] have studied the dynamical properties of quantum entanglement in the integrable Jaynes-Cummings model with Kerr-nonlinearity with various Kerr coupling parameters and initial states, where the initial states are prepared by the coherent states placed in the corresponding phase space described in terms of canonical variables.
The entanglement dynamics after adding Kerr-nonlinearity for IDDJCM are presented in Fig. 8. The addition of nonlinearity $\chi=0.1$ totally destroys the periodic nature of the entanglement for all the subsystem (blue curve in the plots). ESDs from the dynamics are removed for atom A-field a, atom A-atom b and field a-field b subsystems; however, in the case of atom A-atom B, ESDs become longer and peak heights get reduced significantly. As $\chi$ is increased to $0.3$, ESDs get longer in the dynamics of $C(t)$ and also peak which were present till then disappear. For other subsystems, amplitude of $N(t)$ for atom A-field a and filed-field entanglement decrease (green curve); however, for atom A-field b subsystem, $N(t)$ increases. For $\chi=0.7$, which is represented by the red curve in the plots, we observe that $C(t)$ begins to revive again, making the lengths of ESDs shorter. In atom A-field a and field a-field b subsystems, amplitude $N(t)$ decreases significantly, though there is no ESD present in the system. Like atom-atom entanglement, $N(t)$ increases for atom A-field b subsystem and length of ESDs decreases. Though, $N(t)$ decreases at the beginning of the dynamics, it apmplitude increases with time. So, it can be concluded that there is a critical value $\chi$ for which $C(t)$ becomes minimum and beyond that value it increases again. As we increase the nonlinearity further to $\chi=1.0$, same pattern (black curve) is observed as we see for $\chi=0.7$ i.e., the atom-atom and atom A-field b entanglement increase while field-field and atom A-field a entanglement decrease further.
@@
The effects of $\chi$ on entanglement dynamics of different subsystems for SCTS in DJCM are presented in Fig. 9. For $\chi=0.1$, ESD in $C(t)$ increases considerably since the smaller peaks disappear from the dynamics (Fig. 9(a), blue curve). For the other subsystems, the dynamics of $N(t)$ does not change considerably for small values of $\chi$. As $\chi$ is increased to $0.3$, the amplitudes of $C(t)$ and $N(t)$ increase; however, ESDs are not removed from the dynamics. For other entanglements, $N(t)$ for atom A-field a and field a-field b decrease prominently with the increasing nonlinearity in the system. However, $N(t)$ for the atom A-field b subsystem increases and ESDs are removed from the dynamics. On further increasing the value of $\chi$ to $0.7$ and $1.0$, $C(t)$ increases very significantly with greater amplitude and ESDs are removed from the dynamics. Similar behaviour is observed for atom A-field b entanglement dynamics. On the other hand, for atom A-field a and field-field entanglement, $N(t)$ decreases prominently with these increased values. This shows that with the addition of $\chi$ to the system, the entanglement is transferred in the atom-atom and atom A-field b subsystems from the atom A-field b and field a-field b subsystems.
@@ 6 Effects of spin-spin Ising interaction between the two atoms on entanglement dynamics
where $J_{z}$ is the coupling strength between the two atoms. $J_{z}$ has the unit of energy. Recently people have studied the Ising-type interaction for different configurations. Ghoshal et. al., have investigated the entanglement dynamics of the quenched disordered double Jayens-Cummings model in the presence of Ising-type interaction in the Hamiltonian[31]. In another work, Pandit et. al., have also studied the effects of Ising-type interaction on the entanglement dynamics in the DJCM [28]. Ref.[33] has analyzed the effects of Ising-type interaction on entanglement in DJCM for vacuum, coherent and thermal states of radiation fields. Sadiek et. al.,[80] have investigated the time evolution and asymptotic behaviour of entanglement with including the dipole-dipole and Ising-type interactions in the Hamiltonian.
The effects of Ising interaction on entanglement for IDDJCM are shown in Fig. 10. The plots show very different effects as are observed in DJCM for SCS and G-L states. From Fig. 10(a) and 10(b), it can be seen that for $J_{z}=0.1$ and $0.3$, we cannot distinguish the plots from the plots with $J_{z}=0$ for $C(t)$ and $N(t)$ (see Figs. 1 and 10). However, for atom A-field b entanglement, the amplitude of $N(t)$ comes down and ESDs are also present in the dynamics. In the case of field a-field b entanglement, for the above mentioned values of $J_{z}$, we observe that ESDs are removed from the dynamics and amplitude increases. For $J_{z}=0.7$, small peaks appear along with the large peaks in the dynamics of $C(t)$ (red curve) and also the amplitude decreases. For atom A-field a subsystem, the amplitude of $N(t)$ comes down a little and ESDs are also removed in due course of time. In the case of atom A-field b, we do not see any noticeable change in the dynamics. However, for the field a-field b subsystem, the amplitude increases further, and no ESD is observed. For $J_{z}=1.0$, small peaks in the dynamics of $C(t)$ appear earlier in the dynamics and its height increases with time, while the initial peak decreases. As a result, the duration of the ESDs gets smaller. For other subsystems, $N(t)$ follows the dynamics similar to that corresponding to $J_{z}=0.7$, viz., with decreased peak height for atom A-field a entanglement while increased peak height for atom A-field b and field a-field b subsystems.
@@
The effects of Ising interaction on entanglements in DCJM are depicted in Fig. 11. In this case also, we observe that for small values of $J_{z}$ (0.1 and 0.3), $C(t)$ and $N(t)$ for all subsystems do not change considerably. Only field-field entanglement increases slightly for $J_{z}=0.3$. For $J_{z}=0.7$, length of ESD for $C(t)$ increases slightly, as the small bump disappears from the dynamics. The amplitudes of atom-field entanglement do not change noticeably; however, the length of ESD decreases for atom A-field b entanglement. For $J_{z}=1.0$, two small peaks appear in the dynamics of $C(t)$ which makes the length of ESDs smaller. All other entanglement dynamics remain almost unchanged. So, we see that Ising interaction is not much effective to remove sudden death of entanglement for atom-atom subsystem and also from the atom A-field b subsystem.